1. The speed of light in a material is 0.52c. What is the critical angle of a light ray at the interface between the material and a vacuum?
1
Expert's answer
2021-12-01T22:46:55-0500

"{v\,of\,light\,in\,Material=0.52c}\\
{\theta_c=sin^{-1}(\frac{n_{light}}{n_{material}})}\\
{n_{material}=\frac{c}{v}}\\
{n_{material}=\frac{c}{0.52c}}\\
{n_{material}=\frac{1}{0.52}}\\
{n_{material}=1.92}\\
{\theta_c=sin^{-1}(\frac{1}{1.92})}\\
{\theta_c=31.39\degree}\\
{\theta_c=31\degree}\\"